A chair of weight 95.0N lies atop a horizontal floor; the floor is not frictionless. You push on the chair with a force of F = 39.0N directed at an angle of 37.0 degrees below the horizontal and the chair slides along the floor. 1)Using Newton's laws, calculate n, the magnitude of the normal force that the floor exerts on the chair.
Answer #1
Concepts and reason

The concepts that are required to solve this problem are Newton’s second law of motion, equilibrium of forces, resolution of forces, and normal reaction.

Free body diagram is drawn to show all the force acting on the body. The condition for the equilibrium of forces along the vertical direction is to be applied and then equations are obtained to determine the magnitude of the normal reaction.

Fundamentals

The Newton’s second law of motion states that the rate of change of momentum is called the force. Mathematically,

F=ddt(p)F = \frac{d}{{dt}}\left( p \right)

Here, pp is the is the momentum,

Write the expression of momentum,

p=mvp = mv

Here, mm is the mass of the body, and vv is the velocity.

Substitute, mvmv for pp in the expression for Newton’s second law of motion .

F=ddt(mv)=mdvdt=ma\begin{array}{c}\\F = \frac{d}{{dt}}\left( {mv} \right)\\\\ = m\frac{{dv}}{{dt}}\\\\ = ma\\\end{array}

Here, FF is the force, and aa is the acceleration.

The expression of the weight of an object is,

W=mgW = mg

Here, gg is the acceleration due to gravity and mm is mass of the object.

Equilibrium of Forces:

If two or more forces which are acting on the body. The resultant of all forces is zero then this condition is called as equilibrium of forces.

Resolution of forces:

In two dimensional axes, any force at any direction contain a position from x and y axis. Hence, this force can be resolve in two components that is horizontal or x-component and vertical or y-component.

Normal reaction: It is the reaction force of the surface on the body, on which the body is placed upon.

Write the conditions for the equilibrium of forces.

ΣFx=0ΣFy=0\begin{array}{c}\\\Sigma {F_x} = 0\\\\\Sigma {F_y} = 0\\\end{array}

Here, ΣFx\Sigma {F_x} is the sum of all the forces along the horizontal direction, and ΣFy\Sigma {F_y} is the sum of all the forces along the vertical direction.

For a force VV making an angle θ\theta with the positive x-axis, the force resolution to obtain the rectangular components is shown as,

Vsino
> Vcoso

The horizontal component (Vx)\left( {{V_x}} \right) and vertical component (Vy)\left( {{V_y}} \right) is,

Vx=Vcosθ{V_x} = V\cos \theta

Vy=Vsinθ{V_y} = V\sin \theta

(1)

Draw the diagram of the provided arrangement.

NX17370
F = 39 N
W
=95 N

Resolve the force along its components and draw a normal reaction upwards perpendicular to the ground. Draw the free body diagram.

F cos37°
F sin37°
W
= 95 N

Consider the free body diagram of the chair and apply the condition of equilibrium along the vertical direction.

ΣFy=0NWFsinθ=0N=W+Fsinθ\begin{array}{l}\\\Sigma {F_y} = 0\\\\N - W - F\sin \theta = 0\\\\N = W + F\sin \theta \\\end{array}

Substitute 90N90{\rm{ N}} for WW , 39N39{\rm{ N}} for FF and 3737^\circ for θ\theta in the above expression.

N=95.0N+39.0N(sin37.0)=118.47N118.5N\begin{array}{c}\\N = 95.0{\rm{ N}} + 39.0{\rm{ N}}\left( {\sin 37.0^\circ } \right)\\\\ = 118.47{\rm{ N}}\\\\ \simeq {\rm{118}}{\rm{.5}}\;{\rm{N}}\\\end{array}

Ans: Part 1

The magnitude of the normal reaction is 118.5N{\rm{118}}{\rm{.5}}\;{\rm{N}} .